thedukeroi:

Kylee manages a small theme park and she has been analyzing the attendance data. Kylee finds that the number of visitors increases exponentially as the temperature increases, and this situation is represented by the function f(x) = 4x. Kylee also finds a linear equation that models the number of people who leave the park early depending on the change in temperature, and it is represented by g(x) = −x + 5. The graph of the two functions is below. Find the solution to the two functions and explain what the solution represents. (10 points)

Treeko:

@lui0210

thtguy:

To find the solution to the two functions f(x) = 4x and g(x) = -x + 5, we need to solve them simultaneously. To do this, we can set the two functions equal to each other and solve for x: 4x = -x + 5 First, we can add x to both sides of the equation to isolate the x term: 4x + x = -x + x + 5 Simplifying: 5x = 5 Next, we divide both sides of the equation by 5 to solve for x: 5x/5 = 5/5 Simplifying: x = 1

Lui0210:

They said exponentially, so I think it's 4^x not 4x lol

toga:

To find the solution to the two functions, we need to set them equal to each other and solve for x. So, we have: 4x = -x + 5 Adding x to both sides, we get: 5x = 5 Dividing both sides by 5, we get: x = 1 Therefore, the solution to the two functions is x = 1. This solution represents the temperature at which the number of visitors to the theme park is equal to the number of people leaving the park early. In other words, at a temperature of 1 degree (whatever unit of temperature is being used), the park will have an equal number of visitors entering and leaving early due to the change in temperature.